OpenStudy (anonymous):

Write an equation of the line containing the given point an parallel to the given line. (7,-6); 7x-9y=4 The equation of the the line is Y=

OpenStudy (anonymous):

7x-9y=4 => slope = 7/9 the equation of the line will be y=mx+b -6=(7/9)*7+b and then solve for b

OpenStudy (anonymous):

im still confuse with this..

OpenStudy (anonymous):

which step?

OpenStudy (anonymous):

the line is parallel to the given line, so it has the same slope

OpenStudy (anonymous):

how would i write the equation for Y=

OpenStudy (anonymous):

hi, first before we get the parallel line equation,: from this equation 7x-9y=4 what is y=?

OpenStudy (anonymous):

y= 7/9x-4/9

OpenStudy (anonymous):

ok good with this line y= 7/9x-4/9 what is your slope m=?

OpenStudy (anonymous):

hmm seems like you dont like an interaction question and answer,,... :D ok after getting the slope of this line y= 7/9x-4/9 , also get the slope of parallel to this line then form an equation of line from the point (7,-6) with that slope of parallel line.. ..

OpenStudy (anonymous):

@ MauriceLong did you get it ?

OpenStudy (anonymous):

the slope is 4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can u help me solve this problem please

OpenStudy (anonymous):

im sorry the slope of parallel line is the same m1=m2 so you can form another line using this slope m1, using the point(7,-6) and the formula y1-y=m(x-x1)

OpenStudy (anonymous):

ok. i am getting a better understanding on this so far.

OpenStudy (anonymous):

from y= mx + b y= 7/9x - 4/9 what is the slope m?

OpenStudy (anonymous):

yesss just answer what i ask you and you will get it ... :D

OpenStudy (anonymous):

try to compare the formula to the one you answered on y=7/9 x -4/9

OpenStudy (anonymous):

would i have too subtract 7/9-4/9

OpenStudy (anonymous):

by comparing the formula from y= mx + b y= 7/9x - 4/9 what is the slope m? we know that b=4/9, what is slope m?

OpenStudy (anonymous):

no you cant subtract because 7/9 x has an x with it and the 4/9 has no x.. so they are different terms

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if it would be 7/9x - 4/9 x then its ok to subtract but this time they are not.. :D go ahead you can ask me more question what do you want to know

OpenStudy (anonymous):